Last night of the May-SDT-run.

In short: Lots of moonlight -- and bad seeing after a while. High wind speed with gusts >16m/s. Humidity too high after 2am. closed before the night was over. Some Data was taken though.

So: What got observed:

0. Flats in U-spec, R-SLOAN, G-SLOAN, B-BESSEL ( 0 and 180deg each), really good ones...

1. Dwarf Leo-T-g -- FAILED (after repairing the OBfile, first Image was saturated >65k counts. checked with test-exposure: same problem. Pointing off solved the saturation problem, but of course the object could not be observed, straylight/moon? not clear, Leo-T-r was not attempted after that.)

2. GRB070518, seeing 1.0-1.1"

3. GRB070518, seeing varied, some problems with the LBC-software therefore another

4. GRB070518, seeing about 1.0"

5. Lockman VLA U6, seeing 1.3", guiding failed for some time during exp #3

6. Lockman VLA U7, 1.3", again guiding loop opened some times.

7. M92_deepR_F2_5

After that humidity rose rapidly, went above 90%, dewpoint just 1C away. Dome closed. Visually a quite hazy layer around us.

-- PeterBuschkamp - 22 May 2007
Topic revision: r1 - 22 May 2007, PeterBuschkamp
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback